OpenStudy (anonymous):

@dan815 the population of a certain city was 112,000 in 1998, and the observed relative growth rate is 4% per year. (c) in what year will the population reach 200,000

OpenStudy (anonymous):

hello

OpenStudy (dan815):

HELLo MY FRIRND

OpenStudy (dan815):

IAM HERE AT YOUR SERVICE!

OpenStudy (dan815):

LETs GET THIS DONE SHALL WE SIR!

OpenStudy (dan815):

the population of a certain city was 112,000 in 1998, and the observed relative growth rate is 4% per year. (c) in what year will the population reach 200,000

OpenStudy (anonymous):

i need help with more than one problem

OpenStudy (dan815):

City_pop_time_(0)=112000,Year1=1998 Growth_Rate=4/100 City_pop_time_(t)=200,000,Year_Final=Year1-t

OpenStudy (dan815):

solve these system of equations

OpenStudy (dan815):

5 Linear systems of equations

OpenStudy (anonymous):

i have to use logs sir exponential equation heres the equation: 200,000=112,000e^0.04t

OpenStudy (dan815):

okay!

OpenStudy (dan815):

that is fine!

OpenStudy (dan815):

just solve for t

OpenStudy (anonymous):

but im not getting the answer of 2008 how many years?

OpenStudy (dan815):

okay umm

OpenStudy (anonymous):

it starts in 1998

OpenStudy (dan815):

200,000=112,000e^0.04t ln(200,000/112,000 *100/4)=t

OpenStudy (anonymous):

where 100 come from

OpenStudy (dan815):

0.04=4/100

OpenStudy (anonymous):

oh u isolated t

OpenStudy (anonymous):

whats ur answer

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

u there Dan

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

i still need help

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

Dan

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

@swissgirl how u do it

OpenStudy (anonymous):

anybody